Status Report

NASA Space Station On-Orbit Status 8 July 2004

By SpaceRef Editor
July 8, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 8 July 2004

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities. However, Jacques van Oene (
has been violating our copyright by posting these reports, removing some, but not all of our original content in the process. In addition, he has refused to reply to email from SpaceRef or the newsgroup moderator.

All ISS systems continue to function nominally except those noted previously or below.   Day 1314 of permanent human station residency, and Day 2058 since first ISS launch (FGB).

Before “savtrak” (breakfast) and physical exercise, Padalka and Fincke completed another session of the Russian crew health-monitoring program’s medical assessment MO-9/Biochemical Urinalysis.  CDR stowed the hardware afterwards.   [MO-9 is conducted regularly every 30 days (and also before and after EVAs) and is one of five nominal Russian medical tests adopted by NASA for US crewmembers for IMG PHS (Integrated Medical Group/Periodic Health Status) evaluation as part of the “PHS/Without Blood Labs” exam.  The analysis uses the sophisticated in-vitro diagnostic apparatus “Urolux” developed originally for the Mir program.  Afterwards, the data were entered in the medical equipment computer (MEC)’s special IFEP software (In-Flight Examination Program).]

Gennady also conducted his fourth session with the European “Neurocog” experiment (still on board from Pedro Duque’s VC5 “Cervantes” program last October).  Today’s activities again focused on virtual rotation in free floating and fixed position passages while recording EEG (electroencephalogram).  The exercise was videotaped, like the previous one two days ago.  [Assisted by FE Fincke, Padalka first activated the EGE-2 computer, then equipped himself with the Halley head electrodes.  After doing the virtual turns passages in fixed state (subject strapped down) and free-floating in zero-G called for by the Neurocog protocol, he downloaded the EEG data to a PCMCIA memory card for subsequent return to Earth, and dismantled the equipment.]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

The crew conducted the second part of the current ADUM/Advanced Diagnostic Ultrasound in Micro-G protocol, for which Fincke first set up the equipment before becoming the subject for today’s “Scan B” abdominal scanning performed by Gennady as operator, supported by reference pictures uplinked last night.  Afterwards the hardware was deactivated; the scan heads were cleaned and stowed as part of closeout operations.    [After activation of the HRF (Human Research Facility) and the video tape recorder (VTR) by the ground early in the morning, Mike powered up the HRF computer and the ADUM hardware.  The abdominal scans of Fincke, who had to fast for ~6 hrs and drink ~350 ml of clear fluid about one hour prior to scan start, were again medically supported from the ground (POIC, Payload Operations & Integration Center) via privatized video and VOX S/G.  The data were also recorded, and the scanning and post-scan activities were videotaped and still-photographed for downlink.  The ADUM files from yesterday and today were sent down after the activity.]

The Science Officer performed the visual T+2d analysis of the Week 10 water samples, collected on 5/6 from the SRV-K hot tap and the EDV container of the water supply system (SVO-ZV), with the WMK (water monitoring kit).  Subsequently, Mike also did another microbial analysis for inflight coliform bacteria (Escherichia coli) detection on samples collected during Week 4 (T+44d).  He then entered the microbiological data in the medical equipment computer.  [The analyses use incubated MCDs (microbial capture devices), SSK (surface sample kit) slides, and MAS (microbial air sampler) Petri dishes.  If Mike’s analysis showed colony growth above specified limits, he was to take digital documentary images.]

Fincke performed the 30-min. checkout of the MedOps cardiac defibrillator, a periodic routine task that is scheduled as soon as possible from Expedition start and every 60 days thereafter.   [For the checkout, the defib is connected to the 120V outlet, equipped with its battery, today #1008, and then allowed to charge, for about five seconds, to a preset energy level (e.g., 100 joules).  After the button-triggered discharge, a console indicator signals success or failure of the test.  The pacing signal was to be downlinked via S-band for 1.5 min.]

In support of today’s payload ops with EXPRESS rack 5 (ER5), Fincke powered up the rack’s laptop and assisted with its remote activation from the ground.  For the next 10 hrs, POIC/Huntsville tracked the power and thermal resources of the ER during its operations.  This included another run of the SNFM (Serial Network Flow Monitor) for a 3-hr. untended autocapture of LAN-1 science network “bridged” data traffic during the day.  Mike then shut ER5 down again.

Continuing the current round of regular monthly preventive maintenance of RS air ventilation systems that he had started yesterday in the FGB, the CDR today worked an hour on the Service Module (SM)’s air ventilation system to replace its four dust collector filters (PF1-4) and discard the old units.

After this morning’s MO-9 biochemical urinalysis test, FE Fincke set up the equipment for part 2 of the MedOps PHS (periodic health status) assessment, which tomorrow will collect blood samples for analysis with the U.S. PCBA (portable clinical blood analyzer) and the Russian MO-10 “Hemokrit” equipment.   [Preparations included an electronic function test and control analysis on the PCBA, unstowing the AMP (ambulatory medical pack), changing out the batteries of the automatic blood pressure cuff, activating the MEC (medical equipment computer) laptop and opening the IFEP (in-flight examination program) on MEC.  The PHS is performed every 30 days by each crewmember and two weeks before landing or as clinically indicated.]

The CDR completed the routine maintenance of the SM’s SOZh environment control & life support system, which today included the regular weekly inspection of the BRPK air/condensate water separator system, while Mike Fincke prepared the regular IMS “delta” file for IMS database update.

An uplink of new software patches for the Node MDM (multiplexer/demultiplexer, computer) was scheduled for later today.  The upload will be supported by the crew with hard disk “ghosting”.  This will help establishing an early PCS computer capability in the event that the C&C (command & control) MDMs are not available.

Update on EMU troubleshooting:  The crew was commended on yesterday’s “nice work”, during which they ran power usage (amperage) tests on the failed EMU/spacesuits #3013 and #3005.  Since the testing was performed during an LOS (loss-of-signal) period, current data, valve settings and flow data for both suits were downlinked later and are presently being studied.  Preliminary indications are that the water pump in each suit is indeed not working, as originally suspected.   [To verify operation of the impeller in each EMU backpack pump, motor current draw was monitored.  A slightly higher than normal amp draw indicated that the impeller, which is magnetically coupled to the motor, was not turning in both suits.  Other observations, such as the lack of coolant flow, supported the preliminary diagnosis.  More invasive troubleshooting steps are being planned.]

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Update on TsVM reintegration:  Yesterday’s reboot of the Russian segment (RS) Central Computer (TsVM) system to re-integrate all three lanes into the redundant set was successful.  This restart was done ahead of the upgrade restart of TsVM and TVM (Terminal Computer) scheduled for 7/14.
SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities. However, Jacques van Oene (
has been violating our copyright by posting these reports, removing some, but not all of our original content in the process. In addition, he has refused to reply to email from SpaceRef or the newsgroup moderator.

For the upcoming RS software transition to the upgrade version SM 7.02, both MCCs are making preparations for their respective segments, with MCC-Houston intend on maximizing S-band coverage during the transition process.  According to current planning, the software transition will take place from 7/14 to 7/16.

Propellant transfer activities from Progress 14P to the ISS are scheduled for 7/13.  The cargo ship brought 865 kg props to the station.

Departure of 14P then follows on or about 8/2, just before the next Russian spacewalk, EVA-10, now scheduled for 8/3.  Until then, both sides will work hard on updating the existing Flight Rules to the new conditions of a two-man Orlan EVA encountered with EVA-9B.

Today’s CEO (crew earth observations) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Lower Amazon River Basin (weather was predicted to be “as clear as it gets”, with fair-weather cumulus.  To help fill out the crew’s long data acquisition of the lower Amazon and its estuary, images of the north coast of the estuary are still needed.  Shooting obliques looking right for this SW-trending feature.  Obliques can be corrected to document changes coast/island shape and morphology), and Patagonian Glaciers (gap in winter storm trains should allow views of the drier east slopes of the Andes.  Prior efforts have yielded good images of all major glacier tongues.  Images of the smaller glacier tongues are still needed.  Nadir pass).

CEO images can be viewed at these websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 9 crew visit:

U.S. & Russian Segment Status  (as of today, 12:56pm EDT)

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is Off, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is still considered failed).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — n/a; temperature (deg C) — 25.6; ppO2 (mmHg) — 177.0; ppCO2 (mmHg) — 3.1.
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 20.7.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 21.0.
  • Node:  Pressure (mmHg) — 752.8; temperature (deg C) — 22.7 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 754.5; temperature (deg C) — 24.8; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 754.7; temperature (deg C) — 24.2; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • (n/a = data not available)
  • PMA-1:  Shell heater temp (deg C) — n/a
  • PMA-2:  Shell heater temp (deg C) — n/a.

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in directed position (Dual angle/”blind” mode, non solar-tracking, biased for drag reduction).
  • SM batteries:  Battery #4 is in “Cycle” mode; all other batteries (7) are on line in “Partial Charge” mode.  
  • FGB batteries:  Battery #4 is off line; battery #2 is in “Cycle” mode; all other batteries (4) are on line in “Partial Charge” mode.

Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3928 kg (8660 lb) as of 7/1/04;  [SM(552) + FGB(2772) + Progress M(639)].  (Capabilities: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs are on line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04; was replaced 6/30/04).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management, until 6/28, following the EVA.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this morning, 4:43am EDT [= epoch]):

  • Mean altitude — 360.0 km
  • Apogee height — 364.0 km
  • Perigee height — 356.0 km
  • Period — 91.7 min.
  • Inclination (to Equator) —  51.6301 deg
  • Eccentricity — 0.0005954
  • Solar Beta Angle — 66.3 deg
  • Orbits per 24-hr. day — 15.70
  • Mean altitude loss in last 24 hours — 40 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 32164

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.